Х + X Let F(x) = 4 - XO < X < 2 (a) Exp с X Test 2_202005.pdf 5 X (5) WhatsApp ... C:/Users/User/Downloads/Test%202_202005.pdQ2 Given the following heat conduction initial-boundary value problem of a thin homogeneous rod, where ?(?,?) represents the temperature. 9??? = ?? ; 0 < ? < 6; ? > 0; B. C. : ?? (0,?) = 0; ?? (6,?) = 0; ? > 0; I. C. : ?(?, 0) = 12 + 5??? ( ? 6 ?) − 4???(2??); 0 < ? < 6 (a) When ? = 0, what would be the temperature at ? = 3? (Use the initial condition.) (3 marks) (b) Determine whether the boundary conditions in this case is Dirichlet, Neumann, or mixed Dirichlet-Neumann. Then, state the general solution for ?(?,?). (Note: you don’t need to derive the general solution ?(?,?).) (7 marks) (c) Find the particular solution of this initial-boundary value problem making use of your general solution in Q2(b). (10 marks)
Answer #1

In this question to evaluate the first part use initial condition as t=0 given and just put x=3.

now for the second part the conditions are Neumann Conditions and then write the solution.

i have written the derivation but you can also skip this part.

For the last part do the comparing as in u(x,0) there are cosine terms in both the general solution and in the initial conditions.

Que-2 Guinen Where the Answol the following heat corduction inftial boundary value problem pooblem of thin homogeneous roo U1Neuman, g the . (6) Determine wheather Bic is Dirichet mixed dinchlet Neumam ther state general solution Ux,t) Now The Boc in3 They ask solution but we cam you to donot derive denire it as follows: quya 410,t) = Ux16,t) Ut - how by separation of valu4 now x(x). = x(x) x(0) *16) = 0 x() - AX(X) =0 Chavartestic equation :- D=t So There are >o A = P2 D2-1=0 thrill cases å20 again (111) Pto 7 X150) - discard. Aco Let = -p2 P ER Xilx) + p2 x 101) = 0 Characteristic q:D? + P2 = Ipi G coope f g sim-Men e 4 t Tn(t) = now wx,t) Xnit) Init) k ulx, t) = EXn(t) in (e) n 20 -nt t 4 u(x, t) bne E en coo BO (2 nzo Cn bn can 227²127 sios ( 2) - 46729x = Qot a les 19 7) + 8g Co2 957) now Constant Hem Constant >> 12 cofth ro) ie So = 5 10 () a, Lex) a =